A chemist has a 60% solution and a 20% solution he mixes them together to make a 35% solution of 140 mL. how much of the 60% solution was used? how much of the 20% solution?
Answer by mananth(16946)

A chemist has a 60% solution and a 20% solution he mixes them together to make a 35% solution of 140 mL. how much of the 60% solution was used? how much of the 20% solution?
...
60% ------------- x ml
20%---------------(140-x)
Mix 35%------------140 ml
..
60x+20(140-x)=35*140
60x+2800-20x=4900
40x=4900-2800
40x=2100
x= 2100/40
x=52.5 ml of 60%
balance 20% = 87.5
...
m.ananth@hotmail.ca